OpenStudy (winsteria):

FAN AND MEDAL the equation that represents that canned good order is 24x + 64 = 384 x= number of minuets for fruit cans y= number of minuets for vegetable cans explain how to calculate the x- and y- intercepts

OpenStudy (winsteria):

@baru

OpenStudy (anonymous):

is this connected to the other question? O.o haha

OpenStudy (winsteria):

yeah I guess so

OpenStudy (anonymous):

okay so to solve for x we just isolate it form y=mx+b to x=y-b/m

OpenStudy (anonymous):

follow?

OpenStudy (winsteria):

yes I think

OpenStudy (winsteria):

could you help me chose the graph that represents the linear equation 24x + 64y = 384

OpenStudy (winsteria):

@atreyu6s3x6

OpenStudy (winsteria):

@mathmate

OpenStudy (winsteria):

@atreyu6s3x6

OpenStudy (anonymous):

sorry my computer glitched

OpenStudy (winsteria):

it's okay

OpenStudy (anonymous):

okay so we want to put it into slope intercept, its easier, can you do that?

OpenStudy (winsteria):

Idk how to for sure

OpenStudy (anonymous):

okat so we want to isolate y so we: 24x+64y=384 subtract 24x from both sides: 64y=-24x+384 follow?

OpenStudy (winsteria):

okay

OpenStudy (anonymous):

now divide both sides by 64

OpenStudy (winsteria):

64 divided by 64 would be 0 and 64 divided by 384 is 6

OpenStudy (winsteria):

so it's 6 but for the graph would it be the first one because that one has the line that goes through 6

OpenStudy (anonymous):

its actually 1 so we get y=-3/8x+6

OpenStudy (anonymous):

can you graph that?

OpenStudy (winsteria):

I've never really done this before but

OpenStudy (anonymous):

its fine give me a moment :)

OpenStudy (winsteria):

okay

OpenStudy (anonymous):

thats the equation graphed :)

OpenStudy (anonymous):

its actually the first one :)

OpenStudy (winsteria):

oh well I thought that one was wrong because I said that one first lol oh sorry but thank you so much :)

OpenStudy (mathmate):

@winsteria @atreyu6s3x6 If all you need are the intercepts from 24x+64y=384 or from a linear equation in any other form, we first realize that at the x-intercept, y=0, and at the y-intercept, x=0. So to find the x-intercept, we substitute 0 for y, and solve for x, thus 24x+64(0)=384 which means x=384/24=16 To find the y-intercept, we substitute 0 for x, and solve for y, thus 24(0)+64y=384 which means y=384/64=6 So the intercepts are (16,0) and (0,6). The appropriate graph would therefore have an x-intercept of 16, and y-intercept of 6.

OpenStudy (winsteria):

thank you